OpenStudy (anonymous):

A and B are both positive integers. (a^1/2 x b^1/3)^6=432. What is ab?

OpenStudy (loser66):

@terenzreignz

terenzreignz (terenzreignz):

Far more difficult than I'm accustomed to :( \[\Large \left(a^{\frac12}\cdot b^{\frac13}\right)^6=432\]

terenzreignz (terenzreignz):

\[\Large \left.a^{3}\cdot b^{2}\right.=432\]

terenzreignz (terenzreignz):

And then, I suppose, factor 432 into primes.

OpenStudy (anonymous):

Factor 432 and you should get 3^3 x 2^4, and that 2^4 is equal to 4^2 therefore the answer is 12 ?

OpenStudy (raden):

a^1/2 x b^1/3)^6=432 a^3b^2= 16 * 27 a^3b^2= 4^2 * 3^3 identicaly the powers, see the bases you would found the number a and b respectively

OpenStudy (anonymous):

thanks :)

OpenStudy (raden):

welcome :)